Charles Berkle is the manager of Nogain Manufacturing and is interested in doing a cost of quality analysis. The following cost and revenue data are available for the most recent year ended December 31:

Charles Berkle is the manager of Nogain Manufacturing and is interested in doing a cost of quality analysis. The following cost and revenue data are available for the most recent year ended December 31 250,000 Sales revenue 140,000 Cost of goods solo Warranty expense Inspection costs Scrap and rework Product returns due to defects Depreciation expense Machine maintenance expense Wage expense Machine breakdown costs Estimated lost sales due to poor quality 19,000 11,000 7,400 6,000 10,000 2,700 35,000 4,000 5,000 a. Classify each of the above costs into the four quality cost categories and prepare a cost of quality report for Nogain. NOGAIN MANUFACTURING Quality Cost Report For the Year Ended Dec. 31, 20 Prevention costs Machine maintenance Inspections Appraisal costs Internal failure costs External failure costs Total quality costs

b. What percentage of sales revenue is being spent on prevention and appraisal activities? (Round your answer to 1 decimal place.) Costs as a Percentage of Sales c. What percentage of sales revenue is being spent on internal and external failure costs? (Round your answer to 1 decimal place.) Costs as a Percentage of Sales
Answer #1

a Quality control report will be as follows :-

Prevention costs
Machine maintenance expense $2700
Total 2700 $2700
Appraisal costs :
Inspection cost 11000
Total 11000 $11000
Internal failure cost :
Scrap & rework 7400
Machine breakdown costs 4000
Total 11400 $11400
External failure costs :
Warranty expense 19000
Product returns due to defects 6000
Estimated lost sales due to poor quality 5000
Total 30000 $30000
Total quality cost $52400

b . Percentage of sales revenue spent on prevention and appraisal activities = prevention + appraisal / revenue

(2700+11000/250000)x100 = 5.48% or 5.5%

c . Percentage of sales revenue spent on internal and external failure cost = internal + external cost /revenue = (11400+30000/250000)x100 = 16.56%. or 16.6%.